• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name krishniw

    Log entry time 16:49:27 on November 9,2000

    Entry number 51763

    keyword=shift summary

    Day shift 9th Nov 2000.

    We are in controlled access by 8 45.
    Mark Jones and Lubomir went down to change the width of the ADC gates for the calorimeter and also move the high voiltage crate back (at 9.30).

    Ron Gilman went down and enabled the radiator without any problems (separate log antry will be made by RG).

    Bob Micheals noticed that half of the right arm VDC planes are not working (not that it matters to the exp).

    Mark Jones asked the mcc to close the hall, and get readyb for tuning the beam to the hall. (at 3 pm)

    Prediction is beam by 8.00 pm.

    At 3.45 mcc caled, they are going to send us some pulsed
    beam.